i need help solving.

-x+3y+1=2; x=0

Answer by rfer(16322)

-x+3y+1=2
-0+3y+1=2
Zero goes away
Subtract 1 from both sides.
3y+1-1=2-1
3y=1
Divide both sides by 3 to isolate the y.
3y/3=1/3
y=1/3